OpenStudy (anonymous):

Will fan and give Medal! please help –3x + 6y = 6

OpenStudy (anonymous):

one sec, i am figuring it out

OpenStudy (anonymous):

Do you need to know X or Y or both?

OpenStudy (anonymous):

Both

OpenStudy (anonymous):

okie, one sec.

OpenStudy (anonymous):

if you cant figure that on out, what abouy this one? Given f(x) = –3x – 7, what is f(–4)

OpenStudy (keeli_grace):

Okay, y= 1/2 and x= -1, I hope that is what you needed! :)

OpenStudy (anonymous):

@Keeli_Grace, it kind of did.

OpenStudy (anonymous):

i do still need hel on this one though:Given f(x) = –3x – 7, what is f(–4)

OpenStudy (keeli_grace):

Alright, give me a moment. :)

OpenStudy (keeli_grace):

Okay, here is the answer and how you would figure it out: f(x) means what you get if you put x into the function. f(4) means what you get if you put 4 into the function. Given f(x), to calculate f(4) you just change each x in f(x) to a 4 and then work it out. So, f(4) = 3(4) + 7 = 12 + 7 = 19.

OpenStudy (anonymous):

anything else jay?

OpenStudy (anonymous):

@eragondragon666, yea one more:Which is a solution to the equation 4x + y = –6? A. (–2, –2) B. (2, –2) C. (–2, 2) D. (2, –4)

OpenStudy (anonymous):

k just a sec im busy on other Q

OpenStudy (anonymous):

k

OpenStudy (keeli_grace):

x=1 and y=-1 :)

OpenStudy (anonymous):

@Keeli_Grace , thats not one of the options

OpenStudy (keeli_grace):

Really? Hunh... Please forgive me, I will try again.

OpenStudy (keeli_grace):

Oh, I was looking at solutions, you need slope intercept. XD My apologies!

OpenStudy (anonymous):

back oh keeli's got it thx keeli

OpenStudy (anonymous):

thx @Keeli_Grace

OpenStudy (keeli_grace):

Not really, I am still trying. I found the solution, but not the slope-intercept. I am still working on it, so if you get it first, by all means, answer it! :)

OpenStudy (anonymous):

u got it yet

OpenStudy (anonymous):

y int=(0,-1) x int=(2,0)

OpenStudy (anonymous):

slope is 1/2 or o.5

OpenStudy (anonymous):

Your question: 4x +y = -6; would be C. If you have 4 x -2 =-8 +2 =-6 X would be -2, Y would be 2

OpenStudy (anonymous):

Thanks